OpenStudy (matagallina):

What is the exponential function modeled by the following table? x f(x) 2 5 3 9 4 17 f(x) = 2^x f(x) = 2^x + 1 f(x) = 3^x f(x) = 3^x + 1

jhonyy9 (jhonyy9):

what is your opinion ,what will be than you know that 2^2 = 4 and 4+1 = 5 or 3^2 = 9 and 2^3 +1 = 8+1 = 9 and 4^2 = 16 so 4^2 +1 = 17 but 2^4 =16 and 2^4 +1 = 17 do you can deciding what will be correct using these above wrote ?

OpenStudy (magallar23):

f(x) = 2^x + 1